# Bibliography management: natbib package

### v1

@@ 5.1.1 $\Lambda$CDM best fit template function
We show, in Figures 6 and 7, some example contour plots for the hyperparameters used for the crossing functions. We focus on the cases of the Union3 and DES-SN5YR template function, as the two latest, distinct supernova compilations, but the overall shape and behaviour of the contours for Pantheon+ template functions are similar.
@@
As elsewhere, we refer to the values $(C^{n}_{0},C^{n}_{1},...,C^{n}_{n})=(1,0,...,0)$ as the standard hyperparameter values at $n^{th}$ order in crossing functions, since they imply deformations to the template function effectively corresponding to identity. We show results up to $2^{nd}$ order in crossing functions, with the remaining relevant orders for the Union3 and DES-SN5YR template functions shown in Appendix B.
In the case of the Union3 template function, for $C^{0}_{i}$ to $C^{2}_{i}$, there is about $1\sigma$ agreement for the hyperparameters of all datasets with the standard values (see Table 8 for the summary of these results). Conversely, the Pantheon+ template function shows good consistency with its own data, but for the other datasets the hyperparameters are only within $2\sigma$ of the standard values. At 3rd order in Crossing Statistics Union3 and Pantheon+ are both consistent both ways (as template function or data) up to $1\sigma$, but DES-SN5YR data is still more than $2\sigma$ away from the standard values.
To understand this, we examined the contours where the DES-SN5YR best fit template function is used. At 0th and 1st order in Crossing Statistics the data do not prefer any behaviour away from the template function, and the hyperparameters are consistent with standard values to within $1\sigma$. However, for higher orders, as was seen for the template functions from Union3 and Pantheon+, the data only allow the standard hyperparameter values at $2\sigma$ or more. This is true even for the DES-SN5YR data itself, which means that it prefers behaviour that includes deformations away from its own best fit. In contrast, Pantheon+ data manages to agree with the DES-SN5YR best fit template function to within $1\sigma$ of the standard hyperparameter values.
Once 4th order in Crossing Statistics is reached (and DES-SN5YR intrinsic consistency is also attained), all datasets are mutually consistent for the template functions tested. It is notable that the contours of the hyperparameters fit to DES-SN5YR data are significantly larger than for Union3 or Pantheon+ data, above $1^{st}$ order in crossing functions.
It is useful to re-examine the $\Delta\chi^{2}$ in this context. In almost all cases from Union3 and Pantheon+ template functions (except the 1st order deformation of Union3, fit to Pantheon+ data [^11] Again, this implies Union3 $\Lambda$CDM best fit is slightly more difficult to match to the Pantheon+ data, i.e. requires more freedom.), the $\Delta\chi^{2}$ already shows improvement over each dataset’s own best fit at order corresponding to $C^{0}_{i}$. This means that they are not only consistent, but consistent with a good fit.
@@
Thus, DES-SN5YR presents a slight difference to the other datasets. When its best fit is used as its own template function [^12] This is true even when even when $z_{max}$ of $\sim 1.1$ is used in the Chebyshev polynomials, see Section 5.2 for the detailed analysis of this case, and Appendix B, Table 10 for the summary of results., consistency with the standard hyperparameter values is only present at $0^{th}$ and $1^{st}$ order, then not re-attained until $4^{th}$ order. The cases of Union3 and Pantheon+ deformed template functions fit to the DES-SN5YR data prefer non-standard hyperparameter values at lower orders too. This indicates that, until $3^{rd}$ order, there is still some trend, other than $\Lambda$CDM, that is preferred by the data, (but in a way that is seemingly not consistent with the other datasets’ best fit template functions, without deformation [^13] See Section 5.1.2 where we use iteratively-smoothed template functions to explore features of each dataset beyond the concordance model.. At $4^{th}$ order, a lower $\chi^{2}$ may be obtained by values of hyperparameters consistent with standard values to within $1\sigma$, only because the uncertainties are large enough to include the standard hyperparameter values.
@@ 5.1.2 Iterative smoothing template function
In this part of the analysis we examine the affect on mutual consistency when the best fit template functions are replaced by iteratively-smoothed functions which fit better to each dataset and are model-independent. With the maximum redshift of some data exceeding that of DES-SN5YR, and no trivial way to extend an iterative smoothing function beyond the data range, only Union3 and Pantheon+ data are used to produce template functions with iterative smoothing [^14] See Section 5.2 where we consider a different redshift range in all data..
We show some example contour plots for the hyperparameters used by the crossing functions in Figure 8. Once again, we use the same definition of standard hyperparameter values, equivalent to no deformation, and we show only Union3 template function results, up to $2^{nd}$ order. The overall behaviour, which shows a shift in the Union3 contours to center around the standard hyperparameter values, is reproduced in the contour plots for the other template function, which we do not show here. The remainder of the Union3 template function contour plots may be found in Appendix B, along with a summary of the results in Table 9.
It must be borne in mind that the iteratively-smoothed functions from one dataset will be generated in such a way as to match that particular dataset more closely than its own best fit. Consistency between datasets in spite of the resulting features beyond the $\Lambda$CDM model, which arise solely from the data in a model-independent fashion, would imply mutual consistency under some underlying, extended model.
Without much freedom (i.e. $(\mu^{TF}(z)+M_{B})\times C_{0}$), both the Union3 and Pantheon+ template functions $(\mu^{TF}$) are only consistent, to any great degree, with their own data. Once the freedom of at least $C^{1}_{i}$ is allowed, then the Pantheon+ template function is able to be consistent with standard hyperparameter values when fit to Union3 and Pantheon+ data.
In contrast, the Union3 template function requires at least $2^{nd}$ order in crossing functions for consistency with Union3 and Pantheon+ data. This would seem to imply that the Pantheon+ template function is slightly easier to make consistent than the Union3 template function [^15] This is similar to what we see in the best fit template function case., but it should be noted that both require up to order 4 in Crossing Statistics before they can achieve a $\Delta\chi^{2}$ larger than the other dataset’s own iteratively-smoothed function [^16] Though the Pantheon+ template is actually fairly comparable already by second order.
@@
These results so far seem to indicate that DES-SN5YR data do not exhibit as high a level of agreement with the undeformed template functions of the other datasets, as compared to the data from Union3 and Pantheon+. However, as seen in the case of the $\Lambda$CDM model, there still exist regions in the space of hyperparameters where all the data are consistent to within at least $2\sigma$, which indicates that particular deformations of an iterative smoothing extension beyond $\Lambda$CDM can result in realisations compatible with each dataset simultaneously.
One final thing to consider, independent of the type of template function used, is that there is also a striking difference in the redshift range that it covers. Since the template functions are informed by the data, either in terms of the best fit model parameters, or the iteratively-smoothed function, it is reasonable to be suspicious of the effect which this has on the consistency test, and whether the difference in redshift range is leading to a spurious measure of inconsistency. To test this effect directly, we performed the analysis a second time, only considering data within the smallest redshift range of the three datasets (i.e. that of DES-SN5YR). The results of this are included in Section 5.2. While this allows the contours of DES-SN5YR hyperparameters to shrink to sizes comparable with the other two datasets, the same behaviours described above are still seen to a large degree. This indicates that the results showing a preference tending away from the standard hyperparameter values, but still with non-negligible overlap, are not merely a result of the difference in the redshift ranges of the data.
@@ 5.2 Low $z_{max}$ limit
In this analysis, we redo the consistency check in an attempt to remove any possible effects of the mismatch in redshift ranges covered by the different datasets. We make an upper redshift cut at $z_{max}\sim 1.1$ corresponding to the maximum redshift in the DES-SN5YR dataset.
@@
Another important change in this analysis is that the $z_{max}$ used in the Chebyshev polynomials is now that of DES-SN5YR ($z_{max}\sim 1.1$). This means that we also test whether the inconsistency found previously was due, in part, to the fact that DES-SN5YR data only had constraining power over a fraction of the range spanned by the Chebyshev polynomials. This is most likely what leads to the effective reduction in the contour size compared to the initial analysis.
@@ 5.2.1 $\Lambda$CDM best fit template function
The only practical difference in the latter case is that instead of deforming the previous best fit template functions from each dataset to fit the DES-SN5YR data covering a portion (about half) of the rescaled $\tilde{z}$ range, the deformations are fixed by the same data spread over the entire range of $\tilde{z}\in[-1,1]$. This means that deformations of a particular order now occur at a smaller scale, relative to the extent of the data. The hyperparameter values thus no longer have to be as far from standard values to fit better to the data (as compared to when the data only covered a part of the $\tilde{z}$ range).
The change in the range of the data relative to the effective Chebyshev polynomial range is also the reason why the degeneracy direction of the contours is prone to change, compared to the full redshift analysis. Now that the preferred shape of the deformations is governed by a subset of the original data, rescaled to fit into the interval $[-1,1]$, the behaviour of the contours may change. The exact effect on the contours is difficult to predict, in general.
@@ 6 Conclusion
In the case of the first null hypothesis, relating to mutual consistency given the best fit to the assumed flat $\Lambda$CDM model, we may say that it is rejected only at slightly above $2\sigma$ level. This in the case of DES-SN5YR data and the undeformed best fit template functions from either of the two other datasets; whereas, those datasets show a higher level of mutual consistency to each other’s standard hyperparameter values (i.e. undeformed best fits). However, all 3 datasets share enough common regions within the extended space of hyperparameters that they may all be generated by the same subset of deformations, to within $1-2\sigma$.
@@
It would be interesting to explore the kind of expansion histories suggested by each the datasets, and their possible theoretical explanation and implications, as well as the volume, in the space of hyperparameters, which corresponds to the subset of deformations to the template functions which are preferred by all of the datasets. We leave these studies to the realm of future works.
Overall, within the framework of our methodology, the three datasets show a good level of mutual consistency, though their preferred hyperparameters do not generally coincide with each other’s best fit $\Lambda$CDM models. In the case of a model-independent template function, with better fit to each dataset’s particular features, this remains the case. However, within the implied model space provided by the deformations from crossing functions, there are always regions, with $1-2\sigma$ agreement in the hyperparameter values, to be found. Thus, the datasets can be said to be mutually consistency at this level, though not necessarily within that portion of the hyperparameter space corresponding to the concordance cosmology. This means that it is at least possible that no individual dataset is solely systematically incorrect, and points towards the possibility of interesting future studies where the corresponding cosmological implications can be explored.
@@ A.1 Crossing functions
Aside from the deformed function being a good fit to the data, we must choose a maximum order in Crossing Statistics at which to truncate the series. Here we choose a maximum polynomial order of 4 for the crossing functions, based on the following considerations,
@@
  • Small fluctuations: We consider only small deformations around the chosen template function, at the level of $3\%$.
@@
    Note: Consistency requires that we compare the different datasets at the same order in crossing, so we can only do this at the highest order at which the above criteria are met, i.e. $n=4$.
    There are some cases, particularly at lower orders, where the fits to Union3 and DES-SN5YR data are deformed by slightly more than $2\%$. Taking this into account, and the maximum order $n=4$ from physical arguments, in all cases considered in the end, the magnitude of percentage of deformation relative to the template function never exceeds $2.5\%$.

### v2

@@ 4.2 Iterative Smoothing
It may be useful here to recall the number of data points and free parameters in each crossing model. The fit, or generated, template function is fixed in each test. The number of data points of interest are from the dataset to which the template function is deformed:
  • Pantheon+: 1550 [30]
  • Union3: 22 [31]
  • DES-SN5YR: 1829 (effective number: 1735, due to treatment of contamination) [32].
Each order in crossing has a number of free parameters equal to the maximum order, plus the free parameters: $M_{B}$ (translational freedom) and $C_{0}$ (multiplicative bias). Hence, the number of free parameters in fitting a fixed template function to a particular dataset at $n^{\rm th}$ order are: $n+2$, i.e. in the range $2-6$ for the orders considered in our analysis.
@@ 5.1 Full redshift range
In this section, we report the initial results found when considering all three datasets across the full range of each in redshift. We note that the redshift range of the DES-SN5YR dataset is approximately half that of the other two, and thus requires some additional consideration. For each additional crossing hyperparameter, we use a Gaussian prior over the range encompassing their standard values, with 10 on either side, e.g. $C_{0}\in[-10,10]$. This is done to prevent large deviations or unphysical behaviours coming from prior volume effects, assuring that those deviations we see are data-driven. Though this prior might superficially seem the same for each combination of template function and data, the fact that the DES-SN5YR redshift range is smaller means that the effective prior, or the interpretation of the hyperparameters, could be different. The most obvious reason is that the $z_{max}$ used in the Chebyshev polynomials is the overall maximum of all the datasets. This means that deforming the DES-SN5YR template function, in the interval $\tilde{z}\in[-1,1]$ of the Chebyshev polynomials, affect the template function at redshifts not directly informed by the original data.
Vice versa, when fitting the other template functions to the DES-SN5YR dataset, it only has constraining power over a fraction of the range spanned by the Chebyshev polynomials, and so it is reasonable to expect that the allowed values of the hyperparameters will cover a larger range (wider contours) to allow stronger deformations, especially at higher redshifts.
The analysis in this Section is thus most relevant for the comparison of Union3 and Pantheon+ across the full range of available redshifts, though we include DES-SN5YR to have a baseline that will allow us to assess the effect of these differences through further analysis (see Section 5.2).
@@ 5.1.1 $\Lambda$CDM best fit template function
We show, in Figures 6 and 7, some example contour plots for the hyperparameters used for the crossing functions. We focus on the cases of the Union3 and DES-SN5YR template function, as the two latest, distinct supernova compilations, but the overall shape and behaviour of the contours for Pantheon+ template functions are similar. As elsewhere, we refer to the values $(C^{n}_{0},C^{n}_{1},...,C^{n}_{n})=(1,0,...,0)$ as the standard hyperparameter values at $n^{th}$ order in crossing functions, since they imply deformations to the template function effectively corresponding to identity. We show results up to $2^{nd}$ order in crossing functions, with the remaining relevant orders for the Union3 and DES-SN5YR template functions shown in Appendix B.
In the case of the Union3 template function, for $C^{0}_{i}$ to $C^{2}_{i}$, there is about $1\sigma$ agreement for the hyperparameters of all datasets with the standard values (see Table 8 for the summary of these results). Conversely, the Pantheon+ template function shows good compatibility with its own data, but for the other datasets the hyperparameters are only within $2\sigma$ of the standard values. At 3rd order in Crossing Statistics Union3 and Pantheon+ are both compatible both ways (as template function or data) up to $1\sigma$, but DES-SN5YR data is still more than $2\sigma$ away from the standard values.
To understand this, we examined the contours where the DES-SN5YR best fit template function is used. At 0th and 1st order in Crossing Statistics the data do not prefer any behaviour away from the template function, and the hyperparameters are compatible with standard values to within $1\sigma$. However, for higher orders, as was seen for the template functions from Union3 and Pantheon+, the data only allow the standard hyperparameter values at $2\sigma$ or more. This is true even for the DES-SN5YR data itself, which means that it prefers behaviour that includes deformations away from its own best fit. In contrast, Pantheon+ data manages to agree with the DES-SN5YR best fit template function to within $1\sigma$ of the standard hyperparameter values.
Once 4th order in Crossing Statistics is reached (and the preferred deformations of the DES-SN5YR template function, when fit to DES-SN5YR data, now include the standard hyperparameters at less than $2\sigma$.), all datasets are mutually compatible for the template functions tested. Note that the contours of the hyperparameters fit to DES-SN5YR data are significantly larger than for Union3 or Pantheon+ data, above $1^{st}$ order in crossing functions.
It is useful to re-examine the $\Delta\chi^{2}$ in this context. In almost all cases from Union3 and Pantheon+ template functions (except the 1st order deformation of Union3, fit to Pantheon+
@@
data [^12] Again, this implies Union3 $\Lambda$CDM best fit is slightly more difficult to match to the Pantheon+ data, i.e. requires more freedom.), the $\Delta\chi^{2}$ already shows improvement over each dataset’s own best fit at order corresponding to $C^{0}_{i}$. This means that they are not only consistent, but also fit the data well.
@@
Thus, DES-SN5YR presents a slight difference to the other datasets. When its best fit is used as its own template function [^13] This is true even when even when $z_{max}$ of $\sim 1.1$ is used in the Chebyshev polynomials, see Section 5.2 for the detailed analysis of this case, and Appendix B, Table 10 for the summary of results., compatibility with the standard hyperparameter values is only present at $0^{th}$ and $1^{st}$ order, then not re-attained until $4^{th}$ order. The cases of Union3 and Pantheon+ deformed template functions fit to the DES-SN5YR data prefer non-standard hyperparameter values at lower orders too. This indicates that, until $3^{rd}$ order, there is still some trend, other than $\Lambda$CDM, that is preferred by the data, (but in a way that is seemingly not compatible with the other datasets’ best fit template functions, without deformation [^14] See Section 5.1.2 where we use iteratively-smoothed template functions to explore features of each dataset beyond the concordance model.. At $4^{th}$ order, a lower $\chi^{2}$ may be obtained by values of hyperparameters compatible with standard values to within $1\sigma$, only because the uncertainties are large enough to include the standard hyperparameter values.
@@ 5.1.2 Iterative smoothing template function
In this part of the analysis we examine the effect on mutual consistency when the best fit template functions are replaced by iteratively-smoothed functions which more closely follow the data from each dataset and are model-independent. With the maximum redshift of some data exceeding that of DES-SN5YR, and no trivial way to extend an iterative smoothing function beyond the data range, only Union3 and Pantheon+ data are used to produce template functions with iterative smoothing [^15] See Section 5.2 where we consider a different redshift range in all data..
We show some example contour plots for the hyperparameters used by the crossing functions in Figure 8. Once again, we use the same definition of standard hyperparameter values, equivalent to no deformation, and we show only Union3 template function results, up to $2^{nd}$ order. The overall behaviour, which shows a shift in the Union3 contours to center around the standard hyperparameter values, is reproduced in the contour plots for the other template function, which we do not show here. The remainder of the Union3 template function contour plots may be found in Appendix B, along with a result summary in Table 9.
It must be borne in mind that the iteratively-smoothed functions from one dataset will be generated in such a way as to match that particular dataset more closely than its own model best fit. Consistency between datasets in spite of the resulting features beyond the $\Lambda$CDM model, which arise solely from the data in a model-independent fashion, would imply mutual consistency under some underlying, extended model.
With 2 free parameters (i.e. $(\mu^{TF}(z)+M_{B})\times C^{0}_{0}$), both the Union3 and Pantheon+ template functions $(\mu^{TF}$) are only compatible, to any great degree, with their own data. With at least 3 free parameters ($M_{B}$, $C^{1}_{i=1,2}$), the Pantheon+ template function is able to be compatible with standard hyperparameter values when fit to Union3 and Pantheon+ data.
In contrast, the Union3 template function requires at least $2^{nd}$ order in crossing functions for compatibility with Union3 and Pantheon+ data. This would seem to imply that the Pantheon+ template function is slightly easier to make compatible than the Union3 template function [^16] This is similar to what we see in the best fit template function case., but it should be noted that both require up to order 4 in Crossing Statistics before they can achieve a $\Delta\chi^{2}$ larger than the other dataset’s own iteratively-smoothed function [^17] Though the Pantheon+ template is actually fairly comparable already by second order.
@@
These results so far seem to indicate that DES-SN5YR data do not exhibit as high a level of agreement with the undeformed template functions of the other datasets, compared to the data from Union3 and Pantheon+. However, as seen in the case of the $\Lambda$CDM model, there still exist regions in the space of hyperparameters where all the data are consistent to within at least $2\sigma$, which indicates that particular deformations of an iterative smoothing extension beyond $\Lambda$CDM can result in realisations compatible with each dataset simultaneously.
One final thing to consider, independent of the type of template function used, is the striking difference in the redshift range that it covers, which we have previously argued might lead to spurious effects, even though seemingly equivalent priors are used. Since the template functions are informed by the data, either in terms of the best fit model parameters, or the iteratively-smoothed function, it is reasonable to be suspicious of the effect which this has on the consistency test, and whether the difference in redshift range is leading to a spurious measure of inconsistency. We thus perform the analysis a second time, only considering data within the smallest redshift range of the three datasets, i.e. that of DES-SN5YR (see Section 5.2). While this allows the contours of DES-SN5YR hyperparameters to shrink to sizes comparable with the other two datasets, the same behaviours described above are still seen to a large degree. This indicates that the results showing a preference tending away from the standard hyperparameter values, but still with non-negligible overlap, are not merely a result of the difference in the redshift ranges of the data.
@@ 5.2 Low $z_{max}$ limit
In this analysis, we redo the consistency check in an attempt to remove any possible effects of the mismatch in redshift ranges covered by the different datasets, in particular, the DES-SN5YR data. We make an upper redshift cut at $z_{max}\sim 1.1$ corresponding to the maximum redshift in the DES-SN5YR dataset.
@@
Another important change in this analysis is that the $z_{max}$ used in the Chebyshev polynomials is now that of DES-SN5YR ($z_{max}\sim 1.1$). This probes whether the inconsistency found previously was due, in part, to the fact that DES-SN5YR data only had constraining power over a fraction of the range spanned by the Chebyshev polynomials. This is most likely what leads to the effective reduction in the contour size compared to the initial analysis.
@@ 5.2.1 $\Lambda$CDM best fit template function
The only practical difference in the latter case is that instead of deforming the previous best fit template functions from each dataset to fit the DES-SN5YR data covering a portion (about half) of the rescaled $\tilde{z}$ range, the deformations are fixed by the same data spread over the entire range of $\tilde{z}\in[-1,1]$. This means that deformations of a particular order now occur at a smaller scale, relative to the extent of the data. The hyperparameter values thus no longer have to be as far from standard values to achieve a model closer to the data (as compared to when the data only covered a part of the $\tilde{z}$ range).
The change in the range of the data relative to the effective Chebyshev polynomial range is also the reason why the degeneracy direction of the contours is prone to change, compared
to the full redshift analysis. Now that the preferred shape of the deformations is governed by a subset of the original data, rescaled to fit into the interval $[-1,1]$, the behaviour of the contours may change. The exact effect on the contours is difficult to predict, in general.
@@ 6 Conclusion
In the case of the first null hypothesis, relating to mutual consistency given the best fit to the assumed flat $\Lambda$CDM model, we may say that it is rejected only at slightly above $2\sigma$ level. This in the case of DES-SN5YR data and the undeformed best fit template functions from either of the two other datasets; whereas, those datasets show a higher level of mutual consistency to each other’s standard hyperparameter values (i.e. undeformed best fits). However, all 3 datasets share enough regions within the extended space of hyperparameters that they may all be generated by the same subset of deformations, to within $1-2\sigma$.
@@
It would be interesting to explore the kind of expansion histories suggested by each the datasets, and their possible theoretical explanation and implications, as well as the volume, in the space of hyperparameters, which corresponds to the subset of deformations to the template functions which are preferred by all of the datasets. This is left to future works.
Overall, within the framework of our methodology, the three datasets show a good level of mutual consistency, though their preferred hyperparameters do not generally coincide with each other’s best fit $\Lambda$CDM models. In the case of a model-independent template function, with closer fit to each dataset’s particular features, this remains the case. However, within the implied model space provided by the deformations from crossing functions, there are always regions, with $1-2\sigma$ agreement in the hyperparameter values, to be found. Thus, the datasets can be said to be mutually consistent at this level, though not necessarily within that portion of the hyperparameter space corresponding to the concordance cosmology. This means that it is at least possible that no individual dataset is solely systematically incorrect, and points towards the possibility of interesting future studies where the corresponding cosmological implications can be explored.
7 Acknowledgements
The authors would like to thank Rodrigo Calderon and Wuhyun Sohn for useful discussions during the preparation of this work.
@@ A.1 Crossing functions
Aside from the deformed function being a good fit to the data, we must choose a maximum order in Crossing Statistics at which to truncate the series. Here we choose a maximum polynomial order of 4 for the crossing functions, based on the following considerations:
@@
Small fluctuations: We consider only small deformations around the chosen template function, at the level of $\sim 3\%$.
@@
Note: Consistency requires that we compare the different datasets at the same order in crossing, so we can only do this at the highest order at which the above criteria are met, i.e. $n=4$. There are some cases, particularly at lower orders, where the fits to Union3 and DES-SN5YR data are deformed by slightly more than $2\%$. Taking this into account, and the maximum order $n=4$ from physical arguments, in all cases finally considered, the magnitude of deformation relative to the template function never exceeds $2.5\%$.
Degrees of Freedom: We are interested in extended models with additional parameters that are actually justifiable. In some limit of increasing of crossing parameters, the deformed function will eventually have enough freedom for all datasets to be consistent, at the expense of inflated contours (see, for example, the lower panels of Figs 14 and 15). While this isn’t necessarily guaranteed to happen within the maximum order allowed by the above criteria, it should be checked. In order that we actually consider some deformed template function to be a good fit to the data, it is necessary that we do not include superfluous parameters, thereby increasing the degrees of freedom without correspondingly achieving an appreciable decrease in the $\chi^{2}$. Otherwise “consistency”, as we define it, could be trivially obtained, by deliberately enlarging the hyperparameter contours. Thus, we look for extended models with only relatively few additional parameters that show a significant improvement in $\chi^{2}$.
If we assume that the template function is the true underlying model of the data, then the $\Delta\chi^{2}$ of the deformations with respect to the template function will itself follow a chi-square distribution with number of degrees of freedom equal to the number of parameters fit to produce the deformed function. [^19] Technically this is only true for a sufficiently large dataset [34]. In the case of Union3, this might require further investigation; however, given the good agreement with Pantheon+, we do not think it necessary in this work. We may then use the cumulative distribution functions (CDFs), assuming a chi-square distribution with the requisite number of degrees of freedom [35], to calculate the probability that a random draw would produce a given $\Delta\chi^{2}$ value.
Figure 13: The cumulative distribution functions (CDFs) for the $\Delta\chi^{2}$ obtained by each deformation of the template function are shown in solid lines. Each panel signifies a different $\Lambda$CDM best fit template function (Left panel: Pantheon+, Middle panel: Union3, Right panel: DES-SN5YR), while the colour convention is used to show which dataset this template function is deformed to fit. In other words, $1-CDF$ in each plot and for each Crossing order corresponds to the cumulative probability of the given data assuming the template function.
In Fig. 13, we see these probabilities for each template function, deformed to each dataset. The curves correspond to Fig. 5, and may be interpreted in terms of the number of $\sigma$ confidence at which the $\Delta\chi^{2}$ of the deformed function is distinct from that of the particular template function, given some data. The feature of interest is the point at which the decrease in $\chi^{2}$, compared to that of the template function, is no longer sufficient to improve the probability, given the number of additional free parameters. In the full redshift range case, we see that adding higher orders of crossing hyperparameters is still useful in at least one set of deformations for each template function, until at most $4^{th}$ order. For this reason we can reasonably consider the results from Union3 and Pantheon+ up to maximum crossing order of $2$ in the full redshift analysis.
In the case of the reduced redshift range, where DES-SN5YR may be fairly considered, our prior criteria already exclude analysis above $2^{nd}$ order in crossing statistics, and the significance of the corresponding $\Delta\chi^{2}$ in this range is just as high, if not more so.